Benfords Law claims that numbers chosen from very large data files tend to have l as the first nonzero digit disproportion
Answer #1

Null and alternative hypotheses

Ho : p = 0.301

Ha : p < 0.301

Test statistic

Z = p^ - p / sqrt[ p^ * (1-p^)/n]

Where p^ = r/n = 107/402

p^ = 0.266

Z = (0.266-0.301)/sqrt[0.266*0.734/402]

Z = -1.59

P-value for Z = -1.59 and left tailed test

Using Z table

P-value = P( Z < -1.59)

p-value = 0.056

Decision rule : if p-value < \alpha we Reject the null hypothesis otherwise we fail to reject the null hypothesis

Our p-value = 0.056 > 0.01

The p-value is greater than the level of significance so the data are not statistically significant. Thus we fail to reject the null hypothesis.